# Higher-dimensional Teter rings via the canonical trace

### v1

@@ 4. Relations between the nearly Gorenstein and Teter properties
(1): Assume that $r(R)>\text{codim}(R)$. Then, by Section  4 (1), we obtain $r(R)=\text{codim}(R)$, a contradiction. Hence $r(R)\leq\text{codim}(R)$.
(2): This follows from Section  4 (2).
(3): Suppose $r_{0}(R)=1$. Then $R$ is pseudo-Gorenstein. Hence, by [miyashita2025pseudo, Corollary 4.7(1)], $R$ is Gorenstein, a contradiction. Therefore $r_{0}(R)\geq 3$, and in this case (2) yields $r_{0}(R)=r(R)=3$. ∎
@@
Section  4 (3) fails when $\dim(R)<2$. More precisely, when $\dim(R)=0$ there exists a standard graded nearly Gorenstein ring $R$ with $\text{codim}(R)=3$ and $r_{0}(R)=1$, yet $r(R)=2$; in particular $R$ is not level. Indeed, set
@@ 5. The Teter property for special constructions of algebras
In this section, as an application of Theorem 3.13, we mainly discuss the Teter property of fiber products, tensor products, Segre products, and certain Veronese subrings of standard graded rings. Throughout this section, we retain Section  1 and use the notation of Section  2.
@@ 5.1. The Teter property of fiber products of standard graded rings
$\dim R=\operatorname{max}\{\dim S,\dim T\},\qquad\operatorname{depth}R=\operatorname{min}\{\operatorname{depth}S,\operatorname{depth}T,1\}.$
In particular, $R$ is Cohen–Macaulay if and only if $S$ and $T$ are Cohen–Macaulay and $\dim S=\dim T\leq 1$.
@@
Let $f_{1},\dots,f_{n}$ and $g_{1},\dots,g_{m}$ be minimal generators of $\omega_{A}$ and $\omega_{B}$ as $R$-modules, respectively, and choose a homogeneous element $x\in[\omega_{R}]_{0}$ with $\pi(x)=1$. Then the above short exact sequence shows that $\{\iota(f_{1}),\dots,\iota(f_{n}),\iota(g_{1}),\dots,\iota(g_{m}),x\}$ is a generating set of $\omega_{R}$ as an $R$-module. If neither $A$ nor $B$ is regular, it follows from (1) that this is a minimal generating set of $\omega_{R}$, and hence $R$ is level if and only if (i) holds.
@@
Next we show that none of the elements $\iota(g_{j})$ can be removed from a minimal generating set. Without loss of generality, we may assume $j=1$. Suppose that $\iota(g_{1})$ lies in the submodule generated by the other generators and $x$. That is, there exist $a,b_{2},\dots,b_{m},c\in R$ such that $\iota(g_{1})=a\,\iota(f_{1})+\sum_{j=2}^{m}b_{j}\,\iota(g_{j})+c\,x$. Applying $\pi$ to both sides, we obtain $0=\pi(\iota(g_{1}))=a\,\pi(\iota(f_{1}))+\sum_{j=2}^{m}b_{j}\,\pi(\iota(g_{j}))+c\,\pi(x)=c$, and hence $c=0$. Thus $\iota(g_{1})=a\,\iota(f_{1})+\sum_{j=2}^{m}b_{j}\,\iota(g_{j})$. Since $\iota$ is injective, this can be rewritten in $\omega_{A}\oplus\omega_{B}$ as $(0,g_{1})=a\,(f_{1},0)+\sum_{j=2}^{m}b_{j}\,(0,g_{j})$. Using the identification $R_{\mathfrak{m}_{R}}\cong A_{\mathfrak{m}_{A}}\times_{k}B_{\mathfrak{m}_{B}}$, we can write $a=(a_{A},a_{B})$ and $b_{j}=(b_{j,A},b_{j,B})$. Then the right-hand side becomes $(a_{A}f_{1},\ \sum_{j=2}^{m}b_{j,B}g_{j})$, and comparing the second components, we obtain $g_{1}=\sum_{j=2}^{m}b_{j,B}g_{j}$. This contradicts the assumption that $\{g_{1},\dots,g_{m}\}$ is a minimal generating set of $\omega_{B}$. Thus, for every $j$, the element $\iota(g_{j})$ cannot be removed from a minimal generating set.
@@
We may assume that $R_{0}$ is an infinite field. In this case, in the proof of [puthenpurakal2025higher, Proposition 5.2], the condition (3) in Section  3 (c) is verified. By Section  3 (c), this is equivalent to saying that $R$ is Teter. ∎
Corollary 5.7.
Assume that $A$ and $B$ are Cohen–Macaulay generically Gorenstein standard graded rings with $\dim(A)=\dim(B)=1$. Furthermore, suppose that at least one of $A$ and $B$ is not regular. Then the following are equivalent:
  • (1)
    $R=A\times_{\mathbb{k}}B$ is Teter;
  • (2)
    AA and $B$ both have minimal multiplicity.
@@
We first show that (1) implies (2). Assume that $R$ is Teter. By Theorem 1.2 it follows that $R$ is level. By § 5.1 (3), we have $a_{A}=a_{B}=0$. Since $A$ and $B$ are one-dimensional standard graded rings, this means that both $A$ and $B$ have minimal multiplicity.
Next we show that (2) implies (1). We may assume that $A_{0}=B_{0}=\mathbb{k}$ is an infinite field. Under this assumption, together with § 5.1, the rings $A$ and $B$ are each either Gorenstein or Teter, and moreover we have $a_{A}=a_{B}=0$ and both $A$ and $B$ are level (see [eisenbud1984linear, Section 4]). In the Gorenstein case, the hypothesis further implies that they are hypersurfaces. Combining this with Theorem 3.13, we see that in either case, hypersurface or Teter, the equalities $r(A)=\text{codim}(A)$ and $r(B)=\text{codim}(B)$ hold, and both $\operatorname{tr}_{A}(\omega_{A})]_{1}$ and $\operatorname{tr}_{B}(\omega_{B})]_{1}$ contain a non-zerodivisor. Therefore, by § 5.1, the ring $R$ is Teter. ∎
5.2. The Teter property of Veronese subrings
The purpose of this subsection is to investigate the Teter property of suitably low-dimensional Veronese subrings. The next statement is well known to experts, but we include a proof for the reader’s convenience.
@@ 5.2. The Teter property of Veronese subrings
Assume $R$ is standard graded and $1\leq\dim R\leq 2$. If $R$ has minimal multiplicity (i.e. $e(R)=\operatorname{embdim}(R)-\dim(R)+1$), then $R^{(k)}$ also has minimal multiplicity and $e(R^{(k)})=k^{\dim(R)-1}e(R)$ for all $k\in\mathbb{Z}_{>0}$. In particular, $R^{(k)}$ is Gorenstein if and only if $k^{\dim(R)-1}e(R)\leq 2$.
@@
Set $d=\dim R\leq 2$. If $R$ has minimal multiplicity and $d\leq 2$, then its Hilbert series is
@@
Assume that $R$ is standard graded and $\dim R\leq 2$. If $R$ has minimal multiplicity and that it is either Teter or Gorenstein, then $R^{(k)}$ is again Teter or Gorenstein for every $k\geq 1$. In particular, if either $\dim(R)=1$ with $e(R)\geq 3$ or $\dim(R)=2$, then $R^{(k)}$ is Teter for every $k\geq 2$.
@@
When $\dim(R)=0$, note that $R=\mathbb{k}\oplus R_{1}$ because $R$ has minimal multiplicity. Thus $R^{(k)}=\mathbb{k}$ is Gorenstein for $k\geq 2$; for $k=1$ the claim is tautological. Thus we may assume $\dim R=1$ or $\dim R=2$.
Since $R$ is either Teter or Gorenstein, it is in particular generically Gorenstein, so $\omega_{R}$ can be identified with a fractional ideal of $R$. By Theorem 3.13 there exists a non-zerodivisor $f\in[\operatorname{tr}(\omega_{R})]_{1}$. By [miyashita2024linear, Theorem 6.1] we have $\operatorname{tr}_{R^{(k)}}(\omega_{R}^{(k)})\supset({\mathfrak{m}}^{k-1}(f))^{(k)}$. In particular, for the degree-one non-zerodivisor $f^{k}$ of $R^{(k)}$ we obtain $f^{k}\in[\operatorname{tr}_{R^{(k)}}(\omega_{R}^{(k)})]_{1}$. Hence, by Section  4 and § 5.2, the Veronese subring $R^{(k)}$ is either Gorenstein or Teter. In particular, if either $\dim(R)=1$ with $e(R)\geq 3$ or $\dim(R)=2$, then for every $k\geq 2$ we have $k^{\dim(R)-1}e(R)\geq 3$, and hence $R^{(k)}$ is not Gorenstein by § 5.2. Therefore $R^{(k)}$ is Teter. ∎
Corollary 5.10.
@@ 6. Teter numerical semigroup rings
In Section  4 and Section  5, we primarily investigated the Teter property of standard graded rings. In this section, we investigate the Teter property of numerical semigroup rings, which serve as typical examples of rings that are not semi-standard graded. Our goal is to completely characterize the Teter property of numerical semigroup rings in terms of numerical semigroups (see Theorem 6.5). Furthermore, using Theorem 6.5, we discuss the relationships with almost Gorenstein and nearly Gorenstein rings and some examples.
@@
    there exists $\gamma\in\mathbb{Z}_{>0}$ such that $t^{\gamma}\omega_{R_{H}}\subsetneq R_{H}$ and $t^{a_{i}}\in t^{\gamma}\omega_{R_{H}}$ for at least n1n-1 indices $i\in[n]$.
@@
    If $T(H)\neq\emptyset$, we set $\gamma_{H}:=\operatorname{min}T(H)$ and $\delta_{H}:=\gamma_{H}-(\operatorname{F}(H)+a_{1})\geq 0$.
In other words, $R_{H}$ is Teter if and only if $T(H)\neq\emptyset$.
@@
$\Leftarrow$ is clear. We show $\Rightarrow$. Assume $t^{a_{i}}\in t^{\gamma}\omega_{R_{H}}$ for $i\in[n]$. Then there exists $i_{0}\in[\operatorname{type}(H)]$ such that $t^{a_{i}}\in t^{\gamma-\alpha_{i_{0}}}R_{H}$. Since $t^{\gamma-\alpha_{i_{0}}}R_{H}=\sum_{h\in H}t^{\gamma-\alpha_{i_{0}}+h}$, there exists $h\in H$ such that $a_{i}=\gamma-\alpha_{i_{0}}+h$. Since $\gamma\in T_{1}(H)$, we must have $\gamma-\alpha_{i}\in H$ for any $i\in[\operatorname{type}(H)]$. If $\gamma-\alpha_{i_{0}}=0$, it contradicts the fact that $\operatorname{F}(H)<\gamma$ from Section  6 (1), so $\gamma-\alpha_{i_{0}}\in H\setminus\{0\}$. Therefore, by the minimality of the generator $a_{i}$, we obtain $h=0$, and thus $a_{i}=\gamma-\alpha_{i_{0}}$. Moreover, if there exist indices $i_{1},i_{2}$ such that $a_{i}=\gamma-\alpha_{i_{1}}=\gamma-\alpha_{i_{2}}$, then $\alpha_{i_{1}}=\alpha_{i_{2}}$. ∎
With the above preparations in place, we characterize the Teter property of numerical semigroup rings in terms of numerical semigroups.
@@
When $H$ is a numerical semigroup with minimal multiplicity. Then the following are equivalent:
@@
All non-Gorenstein numerical semigroup rings with minimal multiplicity and embedding dimension $3$ are Teter.
@@
If $R_{H}$ is Teter, then the assertion holds because, by Theorem 6.5, one of the following conditions is satisfied:
  1. (1)
    The condition for $s=1$ is equivalent to $a_{3}-a_{2}=\frac{\operatorname{F}(H)}{2}$;
  2. (2)
    The condition for $s=2$ is equivalent to $a_{3}-a_{1}=\frac{\operatorname{F}(H)}{2}$;
  3. (3)
    The condition for $s=3$ is equivalent to $a_{2}-a_{1}=\frac{\operatorname{F}(H)}{2}$.
The converse is obtained by reversing the above argument. ∎
@@
We provide examples of numerical semigroup rings having arbitrary codimension as follows.
@@
In our setting, we have $\operatorname{PF}(H)=\{saq+id-a\mid(q-1)n+1\leq i\leq qn\}$ (see [numata2014numerical, Theorem 3.1 (2)] or [matthews2004numerical]). By Theorem 6.5, if we set $s=1$, the equality
@@
    If $R_{H}$ is nearly Gorenstein but not almost Gorenstein, then it is Teter;
@@
    When $R_{H}$ is almost Gorenstein, the Hilbert–Burch matrix of $R_{H}$ can be expressed as $A=\begin{pmatrix}x_{1}^{a}&x_{2}^{b}&x_{3}^{c}\\ x_{2}&x_{3}&x_{1}\end{pmatrix}$ where $a,b,c$ is positive integers. In this case, $R_{H}$ is Teter if and only if $\operatorname{min}\{a,b,c\}=1$.
@@
We show (2) first. Since it is a numerical semigroup generated by 3 elements, $H$ is pseudo-symmetric. Therefore, by Section  6, it is sufficient to show the existence of a pair of generators whose difference is $\frac{\operatorname{F}(H)}{2}$. By [herzog2021canonical, Theorem 3.3 (3)], $H$ can be represented using $a,b,c>0$ as $H=\langle bc+b+1,ac+c+1,ab+a+1\rangle$ where $(bc+b+1,ac+c+1)=1$. Furthermore, $\operatorname{PF}(H)=\{abc-1,2(abc-1)\}$. By symmetry, we assume $a\leq b\leq c$. (That is, $ab+a+1\leq ac+c+1\leq bc+b+1$). If $a=1$, $R_{H}$ is clearly Teter, so we show the converse. When $R_{H}$ is Teter, one of the following values must equal abc1abc-1:
  1. (a)
    (bc+b+1)(ac+c+1)(bc+b+1)-(ac+c+1);
  2. (b)
    (bc+b+1)(ab+a+1)(bc+b+1)-(ab+a+1);
  3. (c)
    (ac+c+1)(ab+a+1)(ac+c+1)-(ab+a+1);
Assuming equality in (a) or (c) leads to a contradiction with the fact that the embedding dimension of $H$ is 3. Therefore, (b) implies $a=1$. ∎
@@
  1. (a)
    If $\operatorname{embdim}(H)\geq 4$, Section  6 does not hold in general. Indeed, let H=⟨11,12,14,15⟩. Then one can verify that $R_{H}$ is a nearly Gorenstein ring which is not almost Gorenstein and satisfies $\operatorname{type}(H)=\operatorname{embdim}(H)-1$, whereas $R_{H}$ is not Teter because $\operatorname{PF}(H)=\{13,31,32\}$. Moreover, by Theorem 6.5, $R_{H}$ is not Teter.
  2. (b)
    Moreover, even if a numerical semigroup has minimal multiplicity and embedding dimension at least $4$, its numerical semigroup ring need not be Teter. Indeed, let H = ⟨4,7,9,10 ⟩. Then $R_{H}$ has minimal multiplicity and $\operatorname{PF}(H)=\{3,5,6\}$. Thus $R_{H}$ is not Teter by Theorem 6.5.
@@
If $R_{H}$ has minimal multiplicity and $\operatorname{tr}(\omega_{R_{H}})]_{a_{1}}\not=(0)$, then $R_{H}$ is Teter.
@@
Assume that $H$ has minimal multiplicity and that $\operatorname{tr}(\omega_{R_{H}})]_{a_{1}}\neq 0$. Then $a_{1}\in\operatorname{tr}(H)$, so there exist $i\in[n-1]$, $h\in H$ and $z\in\Omega_{H}^{-}$ such that $a_{1}=h-\alpha_{i}+z$, and in particular $a_{1}+\alpha_{i}\in\Omega_{H}^{-}$ for some $i\in[n-1]$. By the description of $\Omega_{H}^{-}$, this means that for some $i\in[n-1]$ we have $a_{1}+\alpha_{i}-\alpha_{j}\in H\quad\text{for all }j\in[n-1]\setminus\{i\}$. Using the characterization of such an index, we may take i=n1i=n-1, and hence $a_{1}+\alpha_{n-1}-\alpha_{j}\in H\quad\text{for all }j\in[n-1]$. All the elements $a_{1}+\alpha_{n-1}-\alpha_{j}$ lie in the Apéry set $\operatorname{Ap}(H,a_{1})$, where $j\in[n-1]$. By the known structure of $\operatorname{Ap}(H,a_{1})$ in this case, we obtain $a_{1}+\alpha_{n-1}-\alpha_{j}=a_{1}+\alpha_{n-1-j}$ for every $j\in[n-2]$. Cancelling $a_{1}$ gives $\alpha_{n-1}=\alpha_{n-1-j}+\alpha_{j}$ for all $j\in[n-2]$. By Section  6, this condition implies that $R_{H}$ is Teter (and hence almost symmetric), as required. ∎
@@
From the proof above, if $H$ has minimal multiplicity, then $\bigl[\operatorname{tr}(\omega_{R_{H}})\bigr]_{a_{1}}\neq 0$ holds if and only if $R_{H}$ is almost Gorenstein. In particular, by [herzog2019trace, Proposition 6.1 and Theorem 6.6], this condition is also equivalent to $R_{H}$ being nearly Gorenstein.

### v2

@@ 4. Relations between the nearly Gorenstein and Teter properties
(1) and (2): Assume that $r(R)>\text{codim}(R)$. Then, by Section  4 (1), we obtain $r(R)=\text{codim}(R)$, a contradiction. Hence $r(R)\leq\text{codim}(R)$. (2) follows from Section  4 (2).
(3): Suppose $r_{\operatorname{min}}(R)=1$. Then $R$ is pseudo-Gorenstein. Hence, by [miyashita2025pseudo, Corollary 4.7(1)], $R$ is Gorenstein, a contradiction. Therefore $r_{\operatorname{min}}(R)\geq 3$, and in this case (2) yields $r_{\operatorname{min}}(R)=r(R)=3$. ∎
@@
Section  4 (3) fails when $\dim(R)<2$. More precisely, when $\dim(R)=0$ there exists a standard graded nearly Gorenstein ring $R$ with $\text{codim}(R)=3$ and $r_{\operatorname{min}}(R)=1$, yet $r(R)=2$; in particular $R$ is not level. Indeed, set
@@ 5. The Teter property for special constructions of algebras
In this section, as an application of Theorem 3.13, we mainly discuss the Teter property of fiber products, tensor products, Segre products, and certain Veronese subalgebras of standard graded rings.
Throughout this section, we retain Section  1 and use the notation of Section  2.
@@ 5.1. The Teter property of fiber products of standard graded rings
$\dim(R)=\operatorname{max}\{\dim(S),\dim(T)\},\qquad\operatorname{depth}(R)=\operatorname{min}\{\operatorname{depth}(S),\operatorname{depth}(T),1\}.$
In particular, $R$ is Cohen–Macaulay if and only if $S$ and $T$ are Cohen–Macaulay and $\dim(S)=\dim(T)\leq 1$.
@@
Let $f_{1},\dots,f_{n}$ and $g_{1},\dots,g_{m}$ be minimal generators of $\omega_{A}$ and $\omega_{B}$ as $R$-modules, respectively, and choose a homogeneous element $x\in[\omega_{R}]_{0}$ with $\pi(x)=1$. Then the above short exact sequence shows that $\{\iota(f_{1}),\dots,\iota(f_{n}),\iota(g_{1}),\dots,\iota(g_{m}),x\}$ is a generating set of $\omega_{R}$ as an $R$-module. If neither $A$ nor $B$ is regular, then by (1) this is a minimal generating set of $\omega_{R}$. Hence $R$ is level if and only if (i) holds.
@@
Next we show that none of the elements $\iota(g_{j})$ can be removed from a minimal generating set. Without loss of generality, we may assume $j=1$. Suppose that there exist $a,b_{2},\dots,b_{m},c\in R$ such that $\iota(g_{1})=a\,\iota(f_{1})+\sum_{j=2}^{m}b_{j}\,\iota(g_{j})+c\,x$. Applying $\pi$ to both sides, we obtain $0=\pi(\iota(g_{1}))=a\,\pi(\iota(f_{1}))+\sum_{j=2}^{m}b_{j}\,\pi(\iota(g_{j}))+c\,\pi(x)=c$, and hence $c=0$. Thus $\iota(g_{1})=a\,\iota(f_{1})+\sum_{j=2}^{m}b_{j}\,\iota(g_{j})$. Since $\iota$ is injective, this can be rewritten in $\omega_{A}\oplus\omega_{B}$ as $(0,g_{1})=a\,(f_{1},0)+\sum_{j=2}^{m}b_{j}\,(0,g_{j})$. Using the identification $R_{\mathfrak{m}_{R}}\cong A_{\mathfrak{m}_{A}}\times_{k}B_{\mathfrak{m}_{B}}$, we can write $a=(a_{A},a_{B})$ and $b_{j}=(b_{j,A},b_{j,B})$. Then the right-hand side becomes $(a_{A}f_{1},\ \sum_{j=2}^{m}b_{j,B}g_{j})$, and comparing the second components, we obtain $g_{1}=\sum_{j=2}^{m}b_{j,B}g_{j}$. This contradicts the assumption that $\{g_{1},\dots,g_{m}\}$ is a minimal generating set of $\omega_{B}$. Thus, for every $j$, the element $\iota(g_{j})$ cannot be removed from a minimal generating set.
@@
We may assume that $R_{0}$ is infinite. In this case, in the proof of [puthenpurakal2025higher, Proposition 5.2], the condition (3) in Section  3 (c) is verified. By Section  3 (c), this is equivalent to saying that $R$ is Teter. ∎
Corollary 5.7 (see Section  1).
Assume that $A$ and $B$ are Cohen–Macaulay generically Gorenstein standard graded rings with $\dim(A)=\dim(B)=1$, not both of which are regular. Then $R$ is Teter if and only if both of $A$ and $B$ have minimal multiplicity.
@@
Assume that $R$ is Teter. By Theorem 1.2 it follows that $R$ is level. By § 5.1 (3), we have $a_{A}=a_{B}=0$. Since $A$ and $B$ are one-dimensional standard graded rings, this means that both $A$ and $B$ have minimal multiplicity.
Next assume that both of $A$ and $B$ have minimal multiplicity. We may assume that $A_{0}=B_{0}=\mathbb{k}$ is an infinite field. Under this assumption, together with § 5.1, the rings $A$ and $B$ are each either Gorenstein or Teter, and moreover we have $a_{A}=a_{B}=0$ and both $A$ and $B$ are level (see [eisenbud1984linear, Section 4]). In the Gorenstein case, the hypothesis further implies that they are hypersurfaces. Combining this with Theorem 3.13, we see that in either case, hypersurface or Teter, the equalities $r(A)=\text{codim}(A)$ and $r(B)=\text{codim}(B)$ hold, and both $\operatorname{tr}_{A}(\omega_{A})]_{1}$ and $\operatorname{tr}_{B}(\omega_{B})]_{1}$ contain a non-zerodivisor. Therefore, by § 5.1, the ring $R$ is Teter. ∎
5.2. The Teter property of Veronese subalgebras
The purpose of this subsection is to investigate the Teter property of suitably low-dimensional Veronese subalgebras. The next statement is well known to experts, but we include a proof for the reader’s convenience.
@@ 5.2. The Teter property of Veronese subalgebras
Assume $R$ is standard graded and $1\leq\dim(R)\leq 2$. If $R$ has minimal multiplicity (i.e. $e(R)=\operatorname{embdim}(R)-\dim(R)+1$), then $R^{(k)}$ also has minimal multiplicity and $e(R^{(k)})=k^{\dim(R)-1}e(R)$ for all $k\in\mathbb{Z}_{>0}$. In particular, $R^{(k)}$ is Gorenstein if and only if $k^{\dim(R)-1}e(R)\leq 2$.
@@
Set $d=\dim(R)\leq 2$. If $R$ has minimal multiplicity and $d\leq 2$, then its Hilbert series is
@@
Assume that $R$ is standard graded and $\dim(R)\leq 2$. If $R$ has minimal multiplicity and that it is either Teter or Gorenstein, then $R^{(k)}$ is again Teter or Gorenstein for every $k\geq 1$. In particular, if either $\dim(R)=1$ with $e(R)\geq 3$ or $\dim(R)=2$, then $R^{(k)}$ is Teter for every $k\geq 2$.
@@
When $\dim(R)=0$, note that $R=\mathbb{k}\oplus R_{1}$ because $R$ has minimal multiplicity. Thus $R^{(k)}=\mathbb{k}$ is Gorenstein for $k\geq 2$; for $k=1$ the claim is tautological. Thus we may assume $\dim(R)=1$ or $\dim(R)=2$.
Since $R$ is either Teter or Gorenstein, it is generically Gorenstein, so $\omega_{R}$ can be identified with a fractional ideal of $R$. By Theorem 3.13 there exists a non-zerodivisor $f\in[\operatorname{tr}(\omega_{R})]_{1}$. By [miyashita2024linear, Theorem 3.4] we have $\operatorname{tr}_{R^{(k)}}(\omega_{R}^{(k)})\supset({\mathfrak{m}}^{k-1}(f))^{(k)}$. In particular, for the degree-one non-zerodivisor $f^{k}$ of $R^{(k)}$ we obtain $f^{k}\in[\operatorname{tr}_{R^{(k)}}(\omega_{R}^{(k)})]_{1}$. Hence, by Section  4 and § 5.2, $R^{(k)}$ is either Gorenstein or Teter. In particular, if either $\dim(R)=1$ with $e(R)\geq 3$ or $\dim(R)=2$, then for every $k\geq 2$ we have $k^{\dim(R)-1}e(R)\geq 3$, and hence $R^{(k)}$ is not Gorenstein by § 5.2. Therefore $R^{(k)}$ is Teter. ∎
Corollary 5.10 (see Section  1).
@@ 6. Teter numerical semigroup rings
In the Section  4 and Section  5, we primarily investigated the Teter property of standard graded rings. In this section, we investigate the Teter property of numerical semigroup rings, which serve as typical examples of rings that are not semi-standard graded. Our goal is to completely characterize the Teter property of numerical semigroup rings in terms of numerical semigroups. Furthermore, using this characterization, we discuss the relationships with almost Gorenstein and nearly Gorenstein rings and some examples.
@@
    $t^{\gamma}\omega_{R_{H}}\subsetneq{R_{H}}$ and for at least n1n-1 indices $i\in[n]$, $t^{a_{i}}\in t^{\gamma}\omega_{R_{H}}$ for some $\gamma\in\mathbb{Z}_{>0}$.
@@
    Let $\gamma_{H}:=\operatorname{min}(T(H))\geq\operatorname{F}(H)+a_{1}$ and $\delta_{H}:=\gamma_{H}-(\operatorname{F}(H)+a_{1})\geq 0$.
In other words, $R_{H}$ is Teter if and only if $T(H)\not$$\emptyset$.
@@
The direction $\Leftarrow$ is clear. We show $\Rightarrow$. Assume $t^{a_{i}}\in t^{\gamma}\omega_{R_{H}}$ for $i\in[n]$. Then there exists $i_{0}\in[\operatorname{type}(H)]$ such that $t^{a_{i}}\in t^{\gamma-\alpha_{i_{0}}}R_{H}$. Since $t^{\gamma-\alpha_{i_{0}}}R_{H}=\sum_{h\in H}t^{\gamma-\alpha_{i_{0}}+h}$, there exists $h\in H$ such that $a_{i}=\gamma-\alpha_{i_{0}}+h$. Since $\gamma\in T_{1}(H)$, we must have $\gamma-\alpha_{i}\in H$ for any $i\in[\operatorname{type}(H)]$. If $\gamma-\alpha_{i_{0}}=0$, it contradicts the fact that $\operatorname{F}(H)<\gamma$ from Section  6 (1), so $\gamma-\alpha_{i_{0}}\in H\setminus\{0\}$. Therefore, by the minimality of the generator $a_{i}$, we obtain $h=0$, and thus $a_{i}=\gamma-\alpha_{i_{0}}$. Moreover, if there exists $\alpha_{i_{1}}$ and $\alpha_{i_{2}}$ which satisfies $a_{i}=\gamma-\alpha_{i_{1}}=\gamma-\alpha_{i_{2}}$, then $\alpha_{i_{1}}=\alpha_{i_{2}}$. Thus such $i_{0}$ is unique. ∎
@@
When $H$ is a numerical semigroup with minimal multiplicity. The following are equivalent:
@@
All numerical semigroup rings with minimal multiplicity and embedding dimension $3$ are Teter.
@@
Assume that $R_{H}$ is Teter. By Theorem 6.5, the condition for $s=1,2,3$ is equivalent to
$a_{3}-a_{2}=\frac{\operatorname{F}(H)}{2},\quad a_{3}-a_{1}=\frac{\operatorname{F}(H)}{2},\quad a_{2}-a_{1}=\frac{\operatorname{F}(H)}{2},$
respectively. Hence the assertion follows. The converse is obtained by reversing the above argument. ∎
@@
Moreover, we provide examples of numerical semigroup rings having arbitrary codimension as follows.
@@
In our setting, we have $\operatorname{PF}(H)=\{saq+id-a:(q-1)n+1\leq i\leq qn\}$ (see [numata2014numerical, Theorem 3.1 (2)] or [matthews2004numerical]). By Theorem 6.5, if we set $s=1$, the equality
@@
    If $R_{H}$ is nearly Gorenstein but not almost Gorenstein, then $R_{H}$ is Teter;
@@
    When $R_{H}$ is almost Gorenstein, the Hilbert-Burch matrix of $R_{H}$ can be expressed as $A=\begin{pmatrix}x_{1}^{a}&x_{2}^{b}&x_{3}^{c}\\ x_{2}&x_{3}&x_{1}\end{pmatrix}$ where a,b,c is positive integers. In this case, $R_{H}$ is Teter if and only if $\operatorname{min}\{a,b,c\}=1$.
@@
We show (2) first. Since it is a numerical semigroup generated by 3 elements, $H$ is pseudo-symmetric. Therefore, by Cor 4, it is sufficient to show the existence of a pair of generators whose difference is $\frac{\operatorname{F}(H)}{2}$. By [herzog2021canonical, Theorem 3.3 (3)], $H$ can be represented using $a,b,c>0$ as $H=\langle bc+b+1,ac+c+1,ab+a+1\rangle$ where $(bc+b+1,ac+c+1)=1$. Furthermore, $\operatorname{PF}(H)=\{abc-1,2(abc-1)\}$. By symmetry, we assume $a\leq b\leq c$. (That is, $ab+a+1\leq ac+c+1\leq bc+b+1$). If $a=1$, $R_{H}$ is clearly Teter, so we show the converse. Assume that $R_{H}$ is Teter. Then abc1abc-1 equals one of bcac+bcbc-ac+b-c, bcab+babc-ab+b-a, and acab+caac-ab+c-a. Assuming equality for (1) or (3) leads to a contradiction with the fact that the embedding dimension of $H$ is 3. Therefore, (2) implies $a=1$. ∎
@@
  1. (1)
    Section  6 generally does not hold when $\operatorname{embdim}(H)\geq 4$. In fact, we can check that $\langle 11,12,14,15\rangle$ is a non-almost Gorenstein nearly Gorenstein ring satisfying $\operatorname{type}(H)=\operatorname{embdim}(H)-1$, but it is not Teter. Actually, this numerical semigroup has $\operatorname{PF}(H)=\{13,31,32\}$ as its pseudo-Frobenius numbers, but by Theorem 6.5, it can not be Teter.
  2. (2)
    Also, even if the semigroup has minimal multiplicity when the embedding dimension is 4 or more, it is not necessarily Teter. Indeed, $H=\langle 4,7,9,10\rangle$ has minimal multiplicity and has $\operatorname{PF}(H)=\{3,5,6\}$ as its pseudo-Frobenius numbers. However, similarly to the above, it follows from Theorem 6.5 that it can not be Teter.
@@
If $H$ has minimal multiplicity and $\operatorname{tr}(\omega_{R_{H}})]_{a_{1}}\not=(0)$, then $R_{H}$ is Teter.
@@
Assume that $H$ has minimal multiplicity and that $\operatorname{tr}(\omega_{R_{H}})]_{a_{1}}\neq 0$. Then $a_{1}\in\operatorname{tr}(H)$, so there exist $i\in[n-1]$, $h\in H$ and $z\in\Omega_{H}^{-}$ such that $a_{1}=h-\alpha_{i}+z$, and in particular $a_{1}+\alpha_{i}\in\Omega_{H}^{-}$ for some $i\in[n-1]$. By the description of $\Omega_{H}^{-}$, this means that for some $i\in[n-1]$ we have $a_{1}+\alpha_{i}-\alpha_{j}\in H\quad\text{for all }j\in[n-1]\setminus\{i\}$. Using the characterization of such an index, we may take i=n1i=n-1, and hence $a_{1}+\alpha_{n-1}-\alpha_{j}\in H\quad\text{for all }j\in[n-1]$. All the elements $a_{1}+\alpha_{n-1}-\alpha_{j}$ lie in the Apéry set $\operatorname{Ap}(H,a_{1})$, where $j\in[n-1]$. By the known structure of $\operatorname{Ap}(H,a_{1})$ in this case, we obtain $a_{1}+\alpha_{n-1}-\alpha_{j}=a_{1}+\alpha_{n-1-j}$ for every $j\in[n-2]$. Cancelling $a_{1}$ gives $\alpha_{n-1}=\alpha_{n-1-j}+\alpha_{j}$ for all $j\in[n-2]$. By Section  6, this condition implies that $H$ is Teter (and hence almost symmetric), as required. ∎
@@
From the proof above, if $H$ has minimal multiplicity, then $\operatorname{tr}(\omega_{R_{H}})]_{a_{1}}\neq 0$ holds if and only if $R_{H}$ is almost Gorenstein.